Physics

A mouse runs along a baseboard in your house. The mouse's position as a function of time is given by x(t)=pt2+qt, with p = 0.36 m/s2and q = -1.40 m/s .

Determine the mouse's average speed between t = 1.0 s and t = 4.0 s.

I have tried everything and the answer is not 0.40 m/s<Homework 1 Item 17 A mouse runs along a baseboard in your house. The mouses position v0.40 m/s î as a function of time is g
Answer #1

x(+) = 2+…+ 9 + finding time when x) = 0 o = P t²+ at its at t=0 and t = -q - P Finding the time when velocity is zero de it

alt) =-1.36 m Now from v=o it starts movena Fowards positive x direction 10 Houm 1036m Also at t= -as its velocity is xeЯo 1.

-0.32m 100 < 0.16m t=48ec t =1.9448ec t=1s Now at t = 4 sec x(4) - 0.36X16 – 1.4x4 __= 0.16___ Now , we know that average speaverage speed = 1.84 averag = 1.84 u-1 =484 = 0.613 m/sec 3